A bus has its maximum number of passengers when it leaves the bus station.
At the first stop, half of the passengers get off. At the next stop 7 people get on and at the next stop 16 people get off.
There are now 17 people on the bus.

Found 2 solutions by longjonsilver, venugopalramana:
Answer by longjonsilver(2297)

Let x = number of passengers at start.

((x/2)+7)-16 = 17
(x/2)+7 = 33
x/2 = 26
x = 26*2
x = 52

jon.

Answer by venugopalramana(3286)

A bus has its maximum number of passengers when it leaves the bus station.
At the first stop, half of the passengers get off. At the next stop 7 people get on and at the next stop 16 people get off.
There are now 17 people on the bus.
LET X BE THE NUMBER IN THE BUS AT THE BEGINING
HALF GET DOWN =X/2...REMAINING = X/2
7 GET IN ...SO TOTAL =X/2 +7
16 GET OFF AT NEXT STOP.
SO PASSENGERS = X/2 +7 - 16 = X/2 -9 = 17
SO X/2 = 17+9=26
X=52....